Can you tell me the absolute value of this equation:

4-8/3x+2/=-20

Answer by stanbon(75887)

Can you tell me the absolute value of this equation:
4-8/3x+2/=-20
-----
-8|3x+2| = -24
----
|3x+2| = 3
-------
3x+2 = 3 or 3x+2 = -3
-------
3x = 1 or 3x = -5
x = 1/3 or x = -5/3
========================
Cheers,
Stan H.